one atom of an element x weighs6.664*10-23gm.the number of gram atoms in 40 kg of it is

Grade:11

4 Answers

Pankaj
9 years ago
Mass of 1 atom = 6.664 x 10-23 gm
1 Gram atom = mass of 1 mole atoms = 6.023x 1023 x mass of one atom
=6.023x 1023 x6.664 x 10-23= 40.13 gm = 0.004 kg
So,
Number of gram atoms in 40 kg = 1000
Qwerty
5 years ago
Answer is 100 Please send me the solution correct but answer is 100 according to FIITJEE  Solution by the experienced teachers 
Amit Seth
5 years ago
WT.of 1mole of atom of X=wt.of 1atom*Av.const
                                             =6.644*10​-23*6.022*1023
                                                  =40g
Thus the at.wt of X=weight in g​/at.wt
                                  =40*1000/40=1000
 
                                           
 
Rishi Sharma
3 years ago
Hello pratap,
The solution is as follows
Given : Mass of 1 atom = 6.664 x 10^-23 gm
1 Gram atom = mass of 1 mole atoms = 6.023x 10^23 x mass of one atom
= 6.023 x 10^23 x 6.664 x 10^-23
= 40.13 gm
= 0.004 kg
So,
Number of gram atoms in 40 kg = 1000
Thanks
I hope above solution will clear your all doubts.
Please feel free to post and ask as much doubts as possible.
All the best.